# Data Mixing Can Induce Phase Transitions in Knowledge Acquisition

### v1

@@ C.1 General Setup
Our experiments use the GPT-NeoX library [Andonian et al., 2023]. For all experiments, we set the batch size as 512 and the sequence length as 2048. For all the experiments in Section 3, we use a Warmup-Stable-Decay (WSD) learning rate schedule with a peak learning rate of $10^{-3}$. We allocate 160 steps for warmup and the final 10% steps for cooldown. We keep other hyperparameters consistent with those used in Pythia.
@@
Let $S$ denote the total number of training tokens. Then, the model sees $rS$ tokens from the knowledge-dense dataset and $(1-r)S$ tokens from the web data. Let $S_{1}$ and $S_{2}$ denote the total sizes (in tokens) of the knowledge-dense dataset and web data. Since $S_{1}$ is small ($<1$B tokens) and $S_{2}$ is large ($>1$T tokens), for the training horizons $S$ considered in our experiments, we typically have $rS>S_{1}$ and $(1-r)S<S_{2}$. In this case, we replicate the knowledge-dense dataset $rS/S_{1}$ times, sample a random $(1-r)S$-token subset from the web data, and then shuffle the combined data.
@@
C.2 Details of Dataset Construction
C.2.1 Constructing the SynBio Dataset
@@
C.2.2 Constructing the WikiBio Dataset
To create the WikiBio dataset, we query Wikidata to gather names and birth dates of individuals from 16 common occupations, identify each person’s biography on Wikipedia by matching the name with the Wikipedia document title, and retain only those biographies where the person’s occupation and birth date are mentioned within the first paragraph. The detailed composition is listed in Table 8. Inspired by Allen-Zhu and Li [2023], we employ Llama-3.1-70B-Instruct to paraphrase the first paragraph of each biography ten times as a form of data augmentation. See Section C.2.2 for the prompt for paraphrasing. An example of the original text and the paraphrased versions are presented below.

### v2

@@ E.1 General Setup
Our experiments use the GPT-NeoX library [gpt-neox-library]. For all experiments, we set the batch size as 512 and the sequence length as 2048. For all the experiments in Section 3, we use a Warmup-Stable-Decay (WSD) learning rate schedule with a peak learning rate of $10^{-3}$. We allocate 160 steps for warmup and the final 10% steps for cooldown. We keep other hyperparameters consistent with those used in Pythia.
@@
Let $S$ denote the total number of training tokens. Then, the model sees $rS$ tokens from the knowledge-dense dataset and (1r)S(1-r)S tokens from the web data. Let $S_{1}$ and $S_{2}$ denote the total sizes (in tokens) of the knowledge-dense dataset and web data. Since $S_{1}$ is small (<1<1B tokens) and $S_{2}$ is large (>1>1T tokens), for the training horizons $S$ considered in our experiments, we typically have $rS>S_{1}$ and $(1-r)S<S_{2}$. In this case, we replicate the knowledge-dense dataset $rS/S_{1}$ times, sample a random (1r)S(1-r)S-token subset from the web data, and then shuffle the combined data.
@@
E.2 Details of Dataset Construction
E.2.1 Constructing the SynBio Dataset
@@
E.2.2 Constructing the WikiBio Dataset
To create the WikiBio dataset, we first query Wikidata to gather names and birth dates of individuals from 16 common occupations. We then identify each person’s Wikipedia page by matching their name with the page title. We retain the first paragraph of each page, as it typically provides a short summary of the person’s life and contains key biographical information. The detailed composition is listed in Table 9. Finally, to align with the evaluation setup in Section 6, we filter the dataset to ensure both the person’s occupation and birth date are explicitly mentioned.
Inspired by allen2023physics, we employ Llama-3.1-70B-Instruct to paraphrase each biography ten times, thereby simulating the real-world scenario where models encounter different variations of the same person’s information during training. See LABEL:tab:paraphrase for the prompt for paraphrasing. An example of the original text and the paraphrased versions are presented below.